# Efficient Bilinear Attention-based Fusion for Medical Visual Question Answering

### v1

@@ IV-B Experimental Setup
We conduct the experiments on a single NVIDIA RTX 3090 (24GB) GPU. The learning rate is set to 0.0005, with a batch size of 32. We train the model for 40 epochs, and the best-performing model on the validation set is saved as the representative model. The Adamax optimizer is used to optimize the model parameters. To minimize the impact of random variation, the OMniBAN model is trained ten times with ten different random seeds, and the average performance and standard deviation is reported in Table II.
At the model level, based on the architecture described in III, we integrate the TCR module [20], which includes a pre-trained question classifier used to distinguish between open and closed questions. By training separate models for each question type, we aim to address the unique characteristics and linguistic structures associated with open and closed questions, enhancing model specialization and accuracy. Table I indicates models that utilize the TCR module with the label ”QCR”.
@@ IV-C Results and Analysis
TABLE I: Comparison of Accuracy (%) on VQA-RAD test set.
Methods Accuracy
@@
MEVF+SAN[5] 40.7 74.1 60.8
MEVF+BAN[5] 43.9 75.1 62.7
MMQ[21] 53.7 75.8 67.0
QCR[20] 60.0 79.3 71.6
PubMedCLIP+BAN(MEVF)[22] 48.6 78.1 66.5
PubMedCLIP+BAN(QCR)[22] 60.1 80.0 72.1
PubMedCLIP+OMniBAN 57.4 80.6 71.4
BiomedCLIP+Transformer[16] 67.6 79.8 75.2
BiomedCLIP+OMniBAN 66.4 80.9 75.1
TABLE II: Ablation Study on VQA-RAD test set (%).
Methods Multi-head Attention Orthogonality Loss Accuracy
@@
MEVF+LSTM+BAN[5] - - 43.9 75.1 62.7
BiomedCLIP+BioBERT+BAN - - 54.3 ± 3.3 77.3 ± 1.9 68.2 ± 2.0
BiomedCLIP+BioBERT+BAN - 64.3 ± 1.1 79.4 ± 0.9 73.4 ± 0.5
BiomedCLIP+BioBERT+BAN 66.4 ± 1.0 80.9 ± 1.3 75.1 ± 0.8
The results of our experiments on the VQA-RAD test set are presented in Table I, comparing different methods across open, closed, and overall accuracy categories. The baseline models MEVF+SAN [5] and MEVF+BAN [5] achieve overall accuracies of 60.8% and 62.7%, respectively. This incremental improvement suggests that BAN has advantages over SAN in capturing cross-modal interactions, likely due to BAN’s bilinear attention structure, which allows for more intricate alignment between image and text features.
The MMQ [21] model further increases overall accuracy to 67.0%, while the QCR [20] model brings a notable improvement, reaching 71.6% overall accuracy. This leap demonstrates the value of integrating question-type classification to specialize the model for handling distinct question types in MedVQA tasks. The PubMedCLIP+BAN (QCR) model attains an accuracy of 72.1%, showcasing how combining BAN with domain-specific encoders like PubMedCLIP provides gains over general-purpose encoders by better capturing the specific semantics of medical images and text.
The integration of OMniBAN with PubMedCLIP and BiomedCLIP produces competitive results in both closed and overall accuracy. For the PubMedCLIP+OMniBAN model, closed question accuracy reaches 80.6%, marking a slight improvement over other PubMedCLIP configurations. The BiomedCLIP+OMniBAN combination achieves 80.9% accuracy on closed questions, slightly outperforming the original Transformer-based BiomedCLIP model on this metric. This performance on closed questions suggests that OMniBAN effectively captures the structured relationships typically associated with closed questions. The overall accuracy of BiomedCLIP+OMniBAN reaches 75.1%, closely matching the Transformer-based BiomedCLIP model at 75.2%, indicating that OMniBAN provides a viable, computationally efficient alternative to Transformer architectures in cross-modal fusion without sacrificing performance.
A noteworthy trend emerges when observing open and closed question performance with OMniBAN integration. In both PubMedCLIP+OMniBAN and BiomedCLIP+OMniBAN models, closed question accuracy sees a slight increase, whereas open question accuracy experiences a minor decline. This result may be attributed to the structured nature of closed questions, which tend to follow specific patterns or have defined answer sets that OMniBAN can capture effectively through its bilinear attention mechanism. Conversely, open questions are more diverse and less constrained in language, potentially benefiting more from the broader contextual capacity of Transformer models. This highlights a nuanced distinction between BAN and Transformer models in handling different question types within MedVQA, suggesting that while BAN is highly effective for structured information, Transformers may still have advantages in dealing with more flexible, open-ended inquiries.
In practical terms, these results underscore the potential of OMniBAN as an alternative to Transformer-based fusion in MedVQA, particularly for applications where computational efficiency is critical. OMniBAN’s performance, closely matching that of Transformer-based BiomedCLIP without the need for extensive pre-training, suggests that BAN-based architectures can provide similar accuracy with reduced computational demands. This is significant for real-world applications, as it enables deployment in resource-limited environments, such as portable devices or edge computing platforms in healthcare settings. Additionally, the integration of domain-specific encoders like BiomedCLIP and PubMedCLIP further enhances BAN’s effectiveness, underlining the importance of specialized feature extraction for domains like medical imaging and question answering.
@@ IV-D Ablation Study
Table II presents an ablation study on the VQA-RAD test set, illustrating how key components of the OMniBAN model contribute to its performance improvements. Starting with the baseline configuration, MEVF+LSTM+BAN [5] achieves 43.9% accuracy on open questions and 75.1% on closed questions, yielding an overall accuracy of 62.7%. This model configuration, which does not include BiomedCLIP or BioBERT, serves as a benchmark for assessing the impact of our modifications.
Replacing the basic image and text encoders with BiomedCLIP and BioBERT leads to a significant performance gain, particularly on open questions (54.3% accuracy), with overall accuracy rising to 68.2%. This highlights the advantage of using domain-specific encoders trained on large-scale biomedical data, allowing the model to capture more relevant medical features from both images and questions. This improvement confirms the value of BiomedCLIP and BioBERT in enhancing baseline performance without additional fusion mechanisms.
Introducing multi-head self-attention further improves model accuracy, raising the overall score to 73.4%. The addition of multi-head self-attention refines intra-modal relationships within each modality, enriching the feature representations before cross-modal fusion. The multi-head structure allows the model to capture diverse perspectives within each input modality, particularly enhancing the model’s ability to process open-ended questions.
Finally, with Orthogonality Loss added, the model achieves an overall accuracy of 75.1%, with noticeable gains in both open and closed question performance. Orthogonality Loss encourages diversity among the glimpses, reducing redundancy and ensuring that each glimpse focuses on unique aspects of the input features. This component proves particularly effective in a multi-modal setting, as it enables the model to leverage complementary information across heads, capturing a broader range of relevant details.
This demonstrates that each component—domain-specific encoders, multi-head self-attention, and Orthogonality Loss—plays a distinct role in improving BAN model. The combination of these components enables the model to achieve levels of accuracy comparable to Transformer-based models, without the need for extensive pre-training. Together, these results highlight the effectiveness of OMniBAN’s architecture in capturing both intra-modal and cross-modal relationships essential for the MedVQA task.
V Conclusion and Future Work
@@
In this paper, we proposed the Orthogonal Multi-head Bilinear Attention Network (OMniBAN) as an efficient fusion approach for Medical Visual Question Answering. Our model integrates a single-layer multi-head self-attention mechanism with bilinear attention networks to efficiently fuse visual and textual features, and add Orthogonality Loss to reduce redundancy among attention glimpses. Compared to traditional BAN and Transformer-based fusion methods, OMniBAN achieves a balanced performance in terms of accuracy and computational efficiency. Specifically, experimental results show that OMniBAN, when combined with BiomedCLIP, slightly outperforms the original Transformer-based BiomedCLIP model in closed-type questions without the need for extensive pre-training. This highlights the ability of OMniBAN to model complex cross-modal interactions effectively and offers a computationally efficient alternative to Transformer-based approaches while maintaining competitive performance.
For future work, there are several directions worth exploring. One potential avenue is to modify the internal attention computation structure of BAN to further enhance its ability to capture cross-modal interactions. Additionally, incorporating more advanced attention mechanisms or integrating external medical knowledge could further improve the model’s performance on MedVQA tasks.

### v3

@@ IV-B Experimental Setup
We conduct our experiments on a single NVIDIA Tesla V100-SXM2 (16GB) GPU. The learning rate is set to 0.0005, and the batch size is 32. The number of heads in Multi-head Attention and glimpses in BAN are set to 8 and 5, respectively. For the orthogonality loss, we adopt a strategy that linearly increases its weight throughout training (up to 0.5). We train the models for 40 epochs on both the VQA-RAD dataset and the SLAKE-EN dataset, and save the best-performing model on the validation set as the representative model. Model parameters are optimized using the Adamax optimizer. To mitigate randomness, we train the OMniBAN model ten times with different random seeds and report the average performance along with the standard deviation in Table III.
At the model level, we integrate the CR approach [24], which uses a pre-trained question classifier to determine whether a given question is open-ended or closed-ended. Based on this classification, the question is routed to one of two specialized sub-models. This design directly addresses the distinct linguistic structures and answer formats in open versus closed questions, which can help the overall model capture the subtle semantic differences of question types more effectively. As shown in Table I and II, models that leverage this approach are marked with “CR”.
@@ IV-C Results and Analysis
TABLE I: Comparison of Accuracy (%) on VQA-RAD[22] test set. The highest accuracy in each column is marked in bold. Results are underlined to indicate the improvement by OMniBAN.
Reference Methods Fusion Methods Accuracy
@@
MEVF[5] SAN 40.7 74.1 60.8
MEVF[5] BAN 43.9 75.1 62.7
MMQ[25] BAN 53.7 75.8 67.0
CR[24] BAN 60.0 79.3 71.6
CPRD[6] BAN 52.5 77.9 67.8
PubMedCLIP(MEVF)[26] BAN 48.6 78.1 66.5
PubMedCLIP(CR)[26] BAN 60.1 80.0 72.1
PubMedCLIP(CR)[26] OMniBAN(Ours) 57.4 80.6 71.4
BiomedCLIP[17] Transformer 67.6 79.8 75.2
BiomedCLIP(CR)[17] OMniBAN(Ours) 66.4 80.9 75.1
TABLE II: Comparison of Accuracy (%) on SLAKE-EN[23] test set. The highest accuracy in each column is marked in bold. Results are underlined to indicate the improvement by OMniBAN.
Reference Methods Fusion Methods Accuracy
@@
MEVF[5] SAN 75.3 78.4 76.5
MEVF[5] BAN 77.8 79.8 78.6
MMQ2[25] BAN - - -
CR[24] BAN 78.8 82.0 80.0
CPRD[6] BAN 79.5 83.4 81.1
PubMedCLIP(MEVF)[26] BAN 76.5 80.4 78.0
PubMedCLIP(CR)[26] BAN 78.4 82.5 80.1
PubMedCLIP(CR)[26] OMniBAN(Ours) 78.1 85.8 81.1
BiomedCLIP[17] Transformer 82.5 89.7 85.4
BiomedCLIP(CR)[17] OMniBAN(Ours) 82.0 89.9 85.1
  • 2MMQ not reported on the SLAKE dataset.
TABLE III: Ablation Study on VQA-RAD and SLAKE-EN test sets (%).
Dataset Image Encoder Text Encoder Multi-head Attention Orthogonality Loss Accuracy
Open Closed All
VQA-RAD BiomedCLIP[17] BioBERT[19] - - 54.3 ± 3.3 77.3 ± 1.9 68.2 ± 2.0
- 64.3 ± 1.1 79.4 ± 0.9 73.4 ± 0.5
66.4 ± 1.0 80.9 ± 1.3 75.1 ± 0.8
SLAKE-EN BiomedCLIP[17] BioBERT[19] - - 79.0 ± 0.6 84.2 ± 1.4 81.1 ± 1.1
- 80.6 ± 0.4 87.2 ± 1.2 83.2 ± 0.5
82.0 ± 0.2 89.9 ± 1.1 85.1 ± 0.6
Comparisons on VQA-RAD Dataset. Table I shows our results on the VQA-RAD test set, comparing open-ended, closed-ended, and overall accuracy. The baseline models MEVF+SAN [5] and MEVF+BAN [5] achieve 60.8% and 62.7% overall accuracy, respectively, with BAN outperforming SAN due to its bilinear attention mechanism. MMQ [25] further raises performance to 67.0%, while CR [24] brings a notable jump to 71.6% by classifying questions into open or closed types before prediction. Leveraging PubMedCLIP with BAN leads to 66.5% accuracy, which increases to 72.1% when combined with CR. Incorporating OMniBAN improves closed-ended performance to 80.6% (PubMedCLIP) and 80.9% (BiomedCLIP), which indicates its strength on questions with restricted answer sets. Although open-ended accuracy dips slightly, overall performance remains competitive, and it demonstrates that OMniBAN can efficiently capture the patterns of closed questions while keeping pace with Transformer-based fusion approaches.
Comparisons on SLAKE-EN Dataset. We observe similar trends on the SLAKE-EN test set in Table II. Again, CR provides a consistent boost over simpler BAN baselines, and OMniBAN excels at closed-ended queries. For instance, PubMedCLIP(CR)+OMniBAN reaches 85.8% on closed-ended questions and 81.1% overall, which improves upon the 82.5% closed-ended accuracy and 80.1% overall accuracy of PubMedCLIP(CR)+BAN. Likewise, BiomedCLIP(CR)+OMniBAN achieves 89.9% closed-ended accuracy, slightly above BiomedCLIP’s 89.7%, but with a marginal decrease in open-ended accuracy.
TABLE IV: Comparison of Computational Efficiency on VQA-RAD [22] Training Set using parameter sizes (M) and FLOPs (M)
Methods Co-Attention OMniBAN
Parameters (M) 31.910 21.659
FLOPs (M) 701.276 182.059
Efficiency Comparison. In order to demonstrate OMniBAN’s high efficiency, we also compare a typical Transformer-based fusion method, i.e., Co-Attention [14] [21] with OMniBAN, as shown in Table IV. The experiment is conducted on the VQA-RAD training set and evaluated via parameter size and FLOPs. To ensure fairness, the Co-Attention fusion approach includes five layers of cross-modal Transformer encoding (alongside one image–text intra-modal attention layer) to match the five glimpses used by OMniBAN’s bilinear attention.
The results turn out that OMniBAN requires fewer parameters (21.659M vs. 31.910M) and significantly fewer FLOPs (182.059M vs. 701.276M), which indicates that its bilinear attention component can effectively reduce computational overhead compared with Transformer-level fusion method. This makes OMniBAN a compelling choice for MedVQA scenarios where efficiency is of priority.
@@ IV-D Ablation Study
We further explore the contribution from each component in the OMniBAN. Table III presents an ablation study on both the VQA-RAD and SLAKE-EN test sets, starting with a baseline model that uses BAN for cross-modal fusion, BiomedCLIP [17] as the image encoder, and BioBERT [19] as the text encoder, without multi-head attention or orthogonality loss. On VQA-RAD, this baseline achieves 54.3% open, 77.3% closed, and 68.2% overall accuracy. Introducing multi-head attention boosts overall accuracy to 73.4%, which reveals the importance of refining intra-modal representations before combining them via BAN. Finally, adding orthogonality loss brings an additional, though smaller, increase to 75.1%. We attribute this limited incremental gain to the inherent constraints of the encoders, which cap the potential benefits of non-overlapping attention glimpses.
On SLAKE-EN, the same pattern emerges. The baseline obtains 79.0% open, 84.2% closed, and 81.1% overall accuracy. Incorporating multi-head attention again yields a notable jump to 83.2% overall, while orthogonality loss provides a further marginal improvement. These findings confirm the value of single-modality attention and orthogonality in enhancing BAN, and suggests that given sufficiently robust features from domain-specific encoders, bilinear fusion can perform competitively with Transformer-based methods for MedVQA.
V Conclusion
@@
In this paper, we introduced the Orthogonal Multi-head Bilinear Attention Network (OMniBAN) as an efficient fusion framework for Medical Visual Question Answering. By combining a single-layer multi-head self-attention mechanism with bilinear attention and employing Orthogonality Loss, OMniBAN balances accuracy and computational efficiency. Experimental results on VQA-RAD and SLAKE-EN show that OMniBAN, when paired with BiomedCLIP, slightly surpasses the original Transformer-based BiomedCLIP model on closed-type questions and achieves quite similar overall accuracy. This indicates that bilinear attention is capable of capturing the structured patterns often found in such questions, and offers a promising alternative to Transformer-based fusion methods without sacrificing performance.
Beyond efficiency, OMniBAN’s underlying design has broader implications for visual–textual interaction and model adaptability. Future research could delve more deeply into these aspects by examining OMniBAN’s robustness under different data distributions or its ability to integrate external medical knowledge. Refinements to BAN’s internal attention mechanisms also present an avenue for further enhancing cross-modal interactions. Nevertheless, our work has a few limitations. We have not evaluated OMniBAN on broader and more specialized MedVQA datasets (e.g., Surgical VQA), and we have only tested two encoders—PubMedCLIP and BiomedCLIP—which leaves the model’s performance with other potential encoders unexamined. These constraints underscore the need for broader experimentation and more diverse ablations. Overall, our findings show that OMniBAN can serve as an effective and efficient choice for Medical Visual Question Answering, which indicates that bilinear attention deserves continued exploration in medical image understanding.